between them, kristen and brad have 68 cousins. brad has 10 more cousins than kristen. how many cousins does kristen have?
a) 42
b) 29
c) 34
d) 39

Answer by user_dude2008(1862)

k+b=68
b=k+10

k+b=68
k+k+10=68
2k+10=68
2k=58
k=29
b=k+10=29+10=39

Answers: k=29, b=39